OpenStudy (anonymous):

Aright triangle has its right angle at (5,7) and the equation of one of its legs is 2x-y-3=0. Find the equation of the other leg.

OpenStudy (anonymous):

help me here @iGreen

OpenStudy (anonymous):

jamaica... i got (3/2, -3)

OpenStudy (anonymous):

your other question was easier i suggest you finish it first it will help you

OpenStudy (anonymous):

i cant do it, please guide me, whats the next step :D

OpenStudy (anonymous):

im only 12 years old, i need to solve this stuff to accelerate to senior high

OpenStudy (anonymous):

on the first x=0 y becomes =13/7 for one point if y=o x becomes =13/3 plot these points draw the line get the slope the question wants you to find the line at right angles thr0 the point (2,1)

OpenStudy (anonymous):

13/7 how did you get this?

OpenStudy (anonymous):

2x-y-3 =0 2(0) - y=3 y= -3/2

OpenStudy (anonymous):

the slope of the line is (-3/7) so one at right angles =(7/3) this is the locus or radius line

OpenStudy (anonymous):

put x=o in the equation you are only left with y=

OpenStudy (anonymous):

2x-y-3 =0 2(0) - y=3 y= -3/2 is this?

OpenStudy (anonymous):

we are working on different equations i am on the first 3x+7y-13=0

OpenStudy (anonymous):

the locus problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, lets solve that first

OpenStudy (anonymous):

i said this was easier and would help with the other

OpenStudy (anonymous):

how did you get (3 , -7) as the slope?

OpenStudy (anonymous):

did you plot the points ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

well can you get the slope?

OpenStudy (anonymous):

i mean 3/-7

OpenStudy (anonymous):

using the slope formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[m = \frac{ 13/7 - 1 }{ 13/3 - 2 }\]

OpenStudy (anonymous):

this?

OpenStudy (anonymous):

your equation 3x+7y-13=0 so 7y =-3x =13 y =3/7(x)-13 so m=-(3/7)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so i will use y-y1 = m(x-x1)?

OpenStudy (anonymous):

the slope you need is at right angles thro (2,1) because that will be the radius line that you use to touch the point (2,1) it could any radius but it will be on this line

OpenStudy (anonymous):



OpenStudy (anonymous):

the intersection is point (2,1)

OpenStudy (anonymous):



OpenStudy (anonymous):

must go think on it